2: Resonance Structures

For each resonance atructureuse curved arrows to show the delocalization of electron pairs that transforms the structure into the next one (to the right) .
Answer #1
Concepts and reason

The resonance structure of a compound is given whose curved arrow mechanism needs to be shown for the generation of the next resonance structure mentioned. Resonance structures are formed by the delocalization of the π\pi electrons and the lone pair of electron either towards the electron deficient species or towards the more electronegative atom.

Fundamentals

• Resonance structure can be drawn by the delocalization of π\pi electron and lone pair of electrons present in a system. Resonance structures are shown by double headed arrow.

The delocalization of π\pi electron and lone pair in a molecule lowers the resonance energy and thus, the molecule is stabilized.

Consider the following example for drawing the resonance structure by delocalization of lone pair and double bond as shown by curved arrows:

ܕܘܝܺ

• Generally, an arrow starts from a nucleophile and ends to an electrophile showing the attack, that is, from a high electron density region to low electron density region. A leaving group leave by taking the electron density.

The direction of arrow basically shows the movement of electrons as shown below:

Η.
Ce
+
CH
electron deficient
species
Electron rich
species

Use the curved arrows to show the conversion of first resonance structure into second without changing any charges as shown below:

(II)

Use the curved arrows to show the conversion of second resonance structure into third without changing any charges as shown below:

NH
(11)

Use the curved arrows to show the conversion of third resonance structure into fourth without changing any charges as shown below:

Use the curved arrows to show the conversion of fourth resonance structure to regenerate the original structure without any formal charges as shown below:

“ΗΝ

Ans:

Hence, the delocalization of electrons to form the given resonance structures can be shown using curved arrows as follows:

NH
(11)
(IV)
Answer #2

Lane bomds delocalization Duning reeman ce pars charpes Occur. Curved arows hrith sructures Reaonance NH2 NH2